# The Mass Distribution of Stars Stripped in Binaries: The Effect of Metallicity

### v1

@@ 2.2 Binary stellar population synthesis
When envelope-stripping is initiated during the main sequence of the donor star (pathway 1 in Fig. 2), the stripped star that is produced should have a slightly lower mass than what we adopt, as the core is not yet fully formed when mass transfer begins (e.g., Yoon et al., 2010). Accounting for this effect would require an expansion of the binary star model grids, which could be a topic of a future study. Here we attempt to quantify the effect on our results (see Appendix A). Approximately 20% of the stripped stars we model go through pathway 1, which is dominant at the high-mass end. Better accounting for main-sequence mass transfer systems would therefore slightly decrease the number of high-mass stripped stars (see Sen et al., 2023). In future studies, the new mapping from Schürmann et al. (2024) could be implemented to better represent the resulting stripped star masses.
@@
In Fig. 3, we use gray circles to plot the initial-to-stripped-star mass relation from Yungelson et al. (2024), who modeled stripped stars at solar metallicity. Their predicted stripped star masses are systematically lower compared to our models. For example, our 10 progenitor results in a stripped star (similar to Yoon et al. 2017), while Yungelson et al. (2024) predict it should result in a stripped star. The initial mass to core mass mapping is relatively unknown and can be affected by convective overshoot, as found by Johnston et al. (2024), who show that convective overshoot can alter the helium core mass by up to 50%. This large uncertainty further emphasizes the importance of finding observable constraints for core mass ratios, which we aim to do with this study.
@@
To obtain stellar properties for the stars in the modeled population, we interpolate properties obtained from the underlying grid of evolutionary models over the logarithm of the stellar mass. We interpolate the properties for main-sequence stars, stripped stars, and post-main sequence stars separately. Because the absolute age of a star is sensitive to its initial mass, for each of these evolutionary phases, we measure the duration of the phase and normalize the star’s evolution by the total duration to reach a smooth interpolation. In this study, we focus on the stripped star mass, the time that envelope-stripping is initiated, and the duration of the stripped star phase. These properties are easy to interpolate over the binary model grids (when ) and can be interpolated from the single star grids by adopting the helium core mass and timescales derived from the main sequence and post-main sequence evolution of the single stars.
@@
We run the code at all four metallicities, adopting a constant star formation rate of , and evaluate the stellar population after 1 Gyr. Because 1 Gyr is approximately the lifetime of a star, the population of stars with initial masses () will be in equilibrium. We therefore focus our population studies on this mass regime.
@@
We choose to explore the mass distribution of stripped stars with stripped star mass (corresponding to , depending on metallicity). As described above, these systems are expected to be in equilibrium. In addition, our binary population synthesis set-up does not correctly account for the evolution of low-mass binaries () and we therefore significantly underpredict the number of low-mass subdwarfs (see however Han et al. 2002, 2003). To capture stochastic variations, we run the code 100 times for each metallicity111Results from the model runs will be available upon publication..
@@ 3.1 The stripped star mass distribution
In Fig. 4, we plot the mass distributions from our population synthesis models for the four metallicities considered, showing how many stripped stars are expected as a function of stripped star mass. The distributions are plotted in red, orange, yellow, and blue, in order of decreasing metallicity. The mass functions displayed in Fig. 4 effectively convolve the initial mass function with mass-dependent effects such as the stellar lifetime and binary fraction, resulting in a different slope than the initial mass function (see also Schneider et al., 2014). We emphasize that these distributions encompass predictions for hot stripped stars (which initiated mass transfer prior to core helium burning). They would exclude partially stripped stars that may result from later interactions (see § 5.1).
For the mass distributions, we use 20 mass bins, logarithmically spaced between 1 and 50 M⊙, corresponding to a bin width of . The solid line in each distribution denotes the mean number of stars in each mass bin over the 100 model runs. The accompanying colored regions denote 1 and 3 standard deviations above and below the mean.
@@ 3.1.1 A power law slope at the low-mass end
Across the four metallicities, we find a negligible difference between the slopes ( and , in order of decreasing metallicity) and the absolute numbers of stripped stars with mass , suggesting that the shape of the intermediate mass regime of the stripped star mass distribution is only marginally affected by metallicity.
@@
The best fit slopes in Fig. 4 are all shallower than that of the input IMF, where we assumed . This suggests that the fraction of stars that become stripped and the larger relative core masses dominate over the longer lifetimes of the lower mass stars.
@@ 3.1.2 A drop at the high-mass end of the low-metallicity mass distributions
In all three sub-solar populations shown in Fig. 4, the distribution departs from the line fitted to the low-mass end, and there is a drop in the mass distribution at a specific threshold in mass, above which very few stripped stars are created through binary interactions. This drop occurs because of the late expansion of massive stars at lower metallicity, causing the Roche lobe in many binary systems not to be filled until during or after helium-core burning has started, resulting in cool or short-lived stripped stars (see Fig. 1 and §1 along with Klencki et al., 2022).
@@ 3.2 Number estimates
Classical WR stars are often thought to be helium stars in the helium core burning phase, potentially stripped through strong stellar winds or binary interaction. Shenar et al. (2020) compared the luminosity of WR stars with evolutionary models and found that helium stars with mass above roughly 7.5, 11, and 17 and in the Milky Way, LMC, and SMC, respectively, should display WR-type spectra (see their Sect. 3.2). For a constant star-formation rate of /yr evaluated after 1 Gyr, we therefore expect binary-stripped stars of WR type in a solar metallicity population, in a population (similar to the LMC), and in a population (similar to the SMC). These numbers are also included in Table 1.
@@ 3.3 A helium star desert
For low stripped star masses, the expected number of stripped stars is largely insensitive to metallicity. For a constant star-formation rate, stripped stars with masses greater than are expected at all metallicities. Similarly, stripped stars with masses greater than are also expected across all metallicities.
However, as already demonstrated in Fig.s 4 and 5, the number of massive stripped stars is highly metallicity dependent. We expect hot binary-stripped stars with mass above for solar metallicity, while not even one is expected for the lowest metallicity (Fig. 6).
For reference, we indicate the predicted mass ranges where the WR phenomenon is observed to occur in the Milky Way, LMC, and SMC in blue hatching in Fig. 6 (Shenar et al., 2020, Table 1). Because the data do not extend to metallicities lower than , the shaded area past this (to the left of ) is an extrapolation. As the figure shows, binary-stripped stars could appear as WR stars at higher metallicities, making them difficult to distinguish from wind-stripped massive stars. However, this becomes increasingly rare with decreasing metallicity, and assuming that helium stars can only be created through envelope-stripping and wind-stripping, we expect a dearth of helium stars with masses and metallicity , since neither binary evolution nor stellar winds should be efficient at producing these (e.g., Georgy et al., 2013; Groh et al., 2019). We label this region of parameter space a “helium star desert.”
@@ 4.1 Predicted numbers of stripped stars
The Milky Way, LMC, and SMC are the closest stellar populations in which we can observe individual stars. These populations have approximately solar (Milky Way), solar (LMC), and solar metallicity (SMC). In the recent past, they have also had relatively constant star-formation rates of 2 /yr (Milky Way, Xiang et al., 2018; Elia et al., 2022), 0.2 /yr (LMC, Harris & Zaritsky, 2009), and 0.05 /yr (SMC, Rubele et al., 2015). Our number estimates for hot, binary-stripped stars can therefore be used to predict the existing number of helium stars in these populations.
We scale our population predictions for , and (see Fig. 5, Table 1 and §3.2) to the star-formation rates of the Milky Way, LMC, and SMC, and thus estimate that they should contain , , and stripped stars with , respectively. Similarly, , , and stripped stars should have sufficient mass to reach core collapse (, Tauris et al., 2015) in the Milky Way, LMC, and SMC, respectively.
@@ 4.2 Comparison with the Milky Way supernova rate
To estimate a rate from our population, we note that the stripped star phase represents about 10% of the total stellar lifetime. In our distribution, the average mass of a stripped star massive enough to explode as a supernova is , corresponding to an initial mass of . We therefore can expect a stripped star of this type to live around 1.5 Myr on average. For a population in equilibrium, this means that we predict the supernova rate from stripped stars to be the number of stripped stars that will explode divided by their average lifetime, which is yr-1 (for a constant star formation rate of /yr). In addition, while there is active debate about whether massive WR stars should explode as core-collapse SN or direct collapse to BHs (Fryer & Kalogera, 2001), we also attempt to account for the contribution of WR stars. Since they are more massive, they only live years. There are WR stars in the Milky Way that are single and therefore not accounted for in our estimate (see §4.4). This means that we need to add yr-1 to the rate of stripped-envelope supernovae, thus reaching 0.0069 yr-1.
@@
The good match between our predicted and the observed stripped-envelope supernova rates provides support for our population predictions and for the mass at which helium stars can reach core collapse.
4.3 Does late expansion actually happen? Is there a drop in the SMC mass distribution?
@@
As described in §3.1.2, our models predict a drop in the SMC mass distribution at because of its low metallicity, which is predicted to cause massive stars to expand late and therefore become stripped late. Scaling by the adopted star-formation rate of the SMC, we expect binary-stripped stars with masses above this drop. By contrast, the solar metallicity model, which does not show a drop in the mass distribution, predicts stripped stars with masses when scaled by the SMC star-formation rate.
@@
The number of stripped stars with in the SMC can therefore be used as a test of the presence of a drop. If the number of such massive stripped stars in the SMC is closer to 20 than 5, it could disprove the theorized late expansion at low metallicity. Alternatively, it could indicate that these stripped stars were created through a different formation scenario than envelope-stripping in binaries (see e.g., § 5.4).
@@
There is already a detected population of massive helium stars in the SMC—the WR stars. Our models predict that envelope-stripping should produce 0.5 WR star in the SMC, assuming that stripped stars with show WR type spectra. However, while the SMC has 12 WR stars (Hainich et al., 2015; Shenar et al., 2016), 5 of them are in such close binary systems ( day orbital periods, Foellmi et al., 2003; Schootemeijer et al., 2024) that they could have been produced by envelope-stripping in binaries (cf. Pols & Marinus, 1994; Shara et al., 2017). Although the numbers are small and stochastic effects play a role, it is intriguing that WR stars are expected for the same star-formation rate but assuming the distribution with no drop. This better match could suggest that stars with metallicity comparable to that of the SMC do not experience late expansion, at least not above . On the other hand, intriguingly, none of the WR binaries in the SMC reside in intermediate period orbits ( days), which would be the likely outcome of binary stripping in models that currently fail to produce stripped stars due to late expansion. To fully test binary models and the phenomenon of late expansion, it would be ideal to map out the complete stripped star distribution in the SMC, at least down to masses below the expected drop at .

### v2

@@ 2.2 Binary stellar population synthesis
While in our population synthesis approach we allowed for the creation of stripped stars through Case A, Case B, and Case B CEE evolution, our binary MESA models simulated only Case B mass transfer, and the mass derived from the single star MESA models is most appropriate for stripped stars created through Case B and Case B CEE evolution. Our use of the same scaling from to for Case A evolution is therefore an approximation and introduces some uncertainty. When envelope-stripping is initiated during the main sequence of the donor star (Case A, pathway 1 in Fig. 2), the stripped star that is produced should have a slightly lower mass than what we adopted, as the core is not yet fully formed when mass transfer begins (e.g., Yoon et al., 2010). Accounting for this effect would require an expansion of the binary star model grids, which could be a topic of a future study. Here we attempt to quantify the effect on our results (see Appendix A). Approximately 20% of the stripped stars we model go through Case A evolution, which is dominant at the high-mass end. Better accounting for main-sequence mass transfer systems would therefore slightly decrease the number of high-mass stripped stars (see Sen et al., 2023). In future studies, the new mapping from Schürmann et al. (2024) could be implemented to better represent the resulting stripped star masses.
@@
In Fig. 3, we use gray circles to plot the initial-to-stripped-star mass relation from Yungelson et al. (2024), who modeled stripped stars at solar metallicity. Their predicted stripped star masses are systematically lower compared to our models. For example, our 10 progenitor results in a stripped star (similar to Yoon et al. 2017), while Yungelson et al. (2024) predict it should result in a stripped star. The initial mass to core mass mapping is relatively unknown and can be affected by convective overshoot, as found by Johnston et al. (2024), who show that convective overshoot can alter the helium core mass by up to 50%. This large uncertainty, further discussed in Sect. 3.1.3, emphasizes the importance of finding observable constraints for core mass ratios, which we aim to do with this study.
@@
To obtain stellar properties for the stars in the modeled population, we interpolated properties obtained from the underlying grid of evolutionary models over the logarithm of the stellar mass. We interpolated the properties for main-sequence stars, stripped stars, and post-main sequence stars separately. Because the absolute age of a star is sensitive to its initial mass, for each of these evolutionary phases, we measured the duration of the phase and normalized the star’s evolution by the total duration to reach a smooth interpolation. In this study, we focused on the stripped star mass, the time that envelope-stripping is initiated, and the duration of the stripped star phase. These properties are easy to interpolate over the binary model grids (when ) and can be interpolated from the single star grids by adopting the helium core mass and timescales derived from the main sequence and post-main sequence evolution of the single stars.
@@
We ran the code at all four metallicities, adopting a constant star formation rate of . This makes it possible to compare our results with an observed population undergoing constant star formation by multiplying by its star formation rate. In our results, we present the number of stripped stars per rather than the fraction of stripped stars. This choice is based on the way our population synthesis code works, as the star formation rate provides a more useful and accurate scaling factor for our results than the total number of stars. We evaluated the stellar population after 1 Gyr. Because 1 Gyr is approximately the lifetime of a star, it is more than sufficient to ensure that the population of stars that is the focus of this work—those with initial masses ()—will be in equilibrium.
@@
We chose to explore the mass distribution of stripped stars with stripped star mass (corresponding to , depending on metallicity). As described above, these systems are expected to be in equilibrium. In addition, our binary population synthesis set-up does not correctly account for the evolution of low-mass binaries () and we therefore significantly underpredict the number of low-mass subdwarfs (see however Han et al. 2002, 2003). To capture stochastic variations, we ran the code 100 times for each metallicity. We ensured that 100 runs is sufficient by running a second set of 100 runs at each metallicity, and we found that the mean number of stripped stars in each mass bin presented in Sect. 3.1 is within for , within for , and for lower masses. 111Results from the model runs are available at https://doi.org/10.5281/zenodo.14811517..
@@ 3.1 The stripped star mass distribution
In Fig. 4, we plot the mass distributions from our population synthesis models for the four metallicities considered, showing how many stripped stars are expected as a function of stripped star mass. The distributions are plotted in red, orange, yellow, and blue, in order of decreasing metallicity. The mass functions displayed in Fig. 4 effectively convolve the initial mass function with mass-dependent effects such as the stellar lifetime and binary fraction, resulting in a different slope than the initial mass function (see also Schneider et al., 2014). We emphasize that these distributions encompass predictions for hot stripped stars (which initiated mass transfer prior to core helium burning). They would exclude partially stripped stars that may result from later interactions (see Sect. 5.1).
For the mass distributions, we use 20 mass bins, logarithmically spaced between 1 and 50 M⊙, corresponding to a bin width of . Because the predicted number of stripped stars per mass bin is normally distributed for all the runs, we use the mean to mark the distribution and the standard deviation to shade possible variations. The solid line in each distribution denotes the mean number of stars in each mass bin over the 100 model runs. The accompanying colored regions denote 1 and 3 standard deviations above and below the mean.
@@ 3.1.1 A power law slope at the low-mass end
Across the four metallicities, we find a negligible (no more than ) difference between the slopes (, and , in order of decreasing metallicity) and the absolute numbers of stripped stars with mass , suggesting that the shape of the intermediate mass regime of the stripped star mass distribution is only marginally affected by metallicity.
@@
The best fit slopes in Fig. 4 are all shallower than that of the input initial mass function, where we assumed . This suggests that the fraction of stars that become stripped and the larger relative core masses dominate over the longer lifetimes of the lower mass stars.
In the solar-metallicity panel, we also display the population predictions from Yungelson et al. (2024), which reaches up to , using a gray dotted line. Notably, their predicted slope is steeper, and they predict significantly fewer stripped stars overall. This is likely due to our different treatments of convective overshoot, which leads to their predicted lower mass stripped stars (see Fig. 3), as well as different initial period distributions and our inclusion of stripped stars formed through the common envelope channel. However, when accounting for these, we find that our results for stripped stars with are in reasonable agreement.
@@ 3.1.2 A drop at the high-mass end of the low-metallicity mass distributions
In all three sub-solar populations shown in Fig. 4, the distribution departs from the line fitted to the low-mass end, and there is a drop in the mass distribution at a specific threshold in mass, above which very few stripped stars are created through binary interactions. This drop occurs because of the late expansion of massive stars at lower metallicity, causing the Roche lobe in many binary systems not to be filled until during or after helium-core burning has started, resulting in cool or short-lived stripped stars (see Fig. 1 and Sect. 1 along with Klencki et al., 2022).
@@
3.1.3 Uncertainties influencing the shape of the mass distribution
Several uncertain physical mechanisms influence the shape of the predicted mass distribution. Here we discuss the primary sources of uncertainty.
In this study, we adopted a relatively high wind mass loss rate for the intermediate and lower mass stripped stars (see Sect. 2.1.1), which could result in an underestimate of the number of lower mass stripped stars. On the other hand, for more massive stripped stars, we used an approximate approach and assumed the stripped star mass to be that of the terminal age main sequence core (see Sect. 2.2 and Fig. 3). Because mass evolution due to wind mass loss was not included, this could result in a moderate overestimate of the number of massive stripped stars.
In our stellar evolution models, we implemented a fixed step-like convective overshoot (see Sect. 2.1.1). Convective boundary mixing is an active field of research, and recent work suggests that the effect increases with stellar mass (e.g., Johnston, 2021; Johnston et al., 2024). If this is true, it would cause the slope of the mass distribution to be shallower than what we predict here, meaning that we overestimate the number of lower mass stripped stars and underestimate the number of higher mass stripped stars. An example of how the mass distribution changes when adopting an overall lower core mass is shown by the dotted line in the top-left panel of Fig. 4, which shows the mass distribution predicted by Yungelson et al. (2024).
The efficiencies of the three evolutionary pathways that we considered (see Sect. 2.2 and Fig. 2) also remain uncertain. Detailed binary evolution models predict that mass transfer can efficiently remove stellar envelopes for the rough parameters that we assume here, but the efficiency of common envelope ejection is substantially less certain (see however, e.g., Dominik et al., 2012). From an observational perspective, the relative efficiencies of these evolutionary channels is largely unconstrained, since a statistical sample of observed stripped stars would be needed to perform suitable constraints. For this reason, we expect some uncertainties in the relative contributions of the three considered evolutionary channels (see also Appendix A).
Observing a population of stripped stars could help constrain these uncertainties. This is one of the goals of the Ultraviolet Explorer mission (UVEX, see Kulkarni et al., 2021), a UV satellite with a planned launch in 2030. In the meantime, we compare our predictions with nearby populations (see Sect. 4) and find reasonable agreement.
@@ 3.2 Number estimates
Our expected number of stripped stars is higher than the predictions from Yungelson et al. (2024) by a factor of a few up to an order of magnitude, depending on mass (see Table 1). We display their predictions for the cumulative number of solar metallicity stripped stars in Fig. 5 using a gray dotted line. Compared to our solar metallicity model, Yungelson et al. (2024) only predict stripped stars with to exist per /yr, while our estimate is closer to .
Classical WR stars are often thought to be helium stars in the helium core burning phase, potentially stripped through strong stellar winds or binary interaction. Shenar et al. (2020) compared the luminosity of WR stars with evolutionary models and found that helium stars with mass above roughly 7.5, 11, and 17 and in the Milky Way, LMC, and SMC, respectively, should display WR-type spectra (see their Sect. 3.2). For a constant star formation rate of /yr evaluated after 1 Gyr, we therefore expect binary-stripped stars of WR type in a solar metallicity population, in a population (similar to the LMC), and in a population (similar to the SMC). These numbers are also included in Table 1.
@@ 3.3 A helium star desert
For low stripped star masses, the expected number of stripped stars is largely insensitive to metallicity. For a constant star formation rate, stripped stars with masses greater than are expected at all metallicities. Similarly, stripped stars with masses greater than are also expected across all metallicities.
However, as already demonstrated in Figs. 4 and 5, the number of massive stripped stars is highly metallicity dependent. We expect hot binary-stripped stars with mass above for solar metallicity, while not even one is expected for the lowest metallicity (Fig. 6).
For reference, we indicate the predicted mass ranges where the WR phenomenon is observed to occur in the Milky Way, LMC, and SMC in blue hatching in Fig. 6 (Shenar et al., 2020, Table 1). Because the data do not extend to metallicities lower than , the shaded area past this (to the left of ) is an extrapolation. As the figure shows, binary-stripped stars could appear as WR stars at higher masses and metallicities, making them difficult to distinguish from the wind-stripped massive stars that are created at these higher masses. However, this becomes increasingly rare with decreasing metallicity, and assuming that helium stars can only be created through envelope-stripping and wind-stripping, we expect a dearth of helium stars with masses and metallicity , since neither binary evolution nor stellar winds should be efficient at producing these (e.g., Georgy et al., 2013; Groh et al., 2019). We label this region of parameter space a “helium star desert.”
@@ 4.1 Predicted numbers of stripped stars
The Milky Way, LMC, and SMC are the closest stellar populations in which we can observe individual stars. These populations have approximately solar (Milky Way), solar (LMC), and solar metallicity (SMC). In the recent past, they have also had relatively constant star formation rates of 2 /yr (Milky Way, Xiang et al., 2018; Elia et al., 2022), 0.2 /yr (LMC, Harris & Zaritsky, 2009), and 0.05 /yr (SMC, Rubele et al., 2015). Our number estimates for hot, binary-stripped stars can therefore be used to predict the existing number of helium stars in these populations.
We scaled our population predictions for , and (see Fig. 5, Table 1 and Sect. 3.2) to the star formation rates of the Milky Way, LMC, and SMC, and thus estimated that they should contain , , and stripped stars with , respectively. Similarly, , , and stripped stars should have sufficient mass to reach core collapse (, Tauris et al., 2015) in the Milky Way, LMC, and SMC, respectively.
@@ 4.2 Comparison with the Milky Way supernova rate
To estimate a rate from our population, we output the duration of the stripped star phase, , for each stripped star as determined from the evolutionary models. We then calculated the stripped envelope supernova (SESN) rate to be the sum of the inverses of the stripped star durations for each star with ; that is,
@@
Scaling according to the Milky Way star formation rate of /yr, we found that . In addition, while there is active debate about whether massive WR stars should explode as core-collapse supernovae or direct collapse to black holes (Fryer & Kalogera, 2001), we also attempted to account for the contribution of WR stars. Since they are more massive, they only live years (Groh et al., 2014). There are WR stars in the Milky Way that are single and therefore not accounted for in our estimate (see Sect. 4.4). This means that we needed to add yr-1 to the rate of stripped-envelope supernovae, thus reaching 0.0090 yr-1.
@@
The reasonable match between our predicted and the observed stripped-envelope supernova rates provides support for our population predictions and for the mass at which helium stars can reach core collapse. However, see Sect. 3.1.3 for the primary sources of uncertainty that could be responsible for the slight mismatch.
4.3 Investigation of late expansion using the SMC mass distribution
@@
As described in Sect. 3.1.2, our models predict a drop in the SMC mass distribution at because of its low metallicity, which is predicted to cause massive stars to expand late and therefore become stripped late. Scaling by the adopted star formation rate of the SMC, we expect binary-stripped stars with masses above this drop. By contrast, the solar metallicity model, which does not show a drop in the mass distribution, predicts stripped stars with masses when scaled by the SMC star formation rate.
@@
The number of stripped stars with in the SMC can therefore be used as a test of the presence of a drop. If the number of such massive stripped stars in the SMC is closer to 20 than 5, it could disprove the theorized late expansion at low metallicity. Alternatively, it could indicate that these stripped stars were created through a different formation scenario than envelope-stripping in binaries (see e.g., Sect. 5.4).
@@
There is already a detected population of massive helium stars in the SMC—the WR stars. Our models predict that envelope-stripping should produce 0.5 WR star in the SMC, assuming that stripped stars with show WR type spectra. However, while the SMC has 12 WR stars (Hainich et al., 2015; Shenar et al., 2016), 5 of them are in such close binary systems ( day orbital periods, Foellmi et al., 2003; Schootemeijer et al., 2024) that they could have been produced by envelope-stripping in binaries (cf. Pols & Marinus, 1994; Shara et al., 2017). Although the numbers are small and stochastic effects play a role, it is intriguing that WR stars are expected for the same star formation rate but assuming the distribution with no drop. This better match could suggest that stars with metallicity comparable to that of the SMC do not experience late expansion, at least not above . On the other hand, intriguingly, none of the WR binaries in the SMC reside in intermediate period orbits ( days), which would be the likely outcome of binary stripping in models that currently fail to produce stripped stars due to late expansion. To fully test binary models and the phenomenon of late expansion, it would be ideal to map out the complete stripped star distribution in the SMC, at least down to masses below the expected drop at .